OpenStudy (ilovecake):

What is the ratio 18 to 27 written as a fraction in lowest terms?

OpenStudy (ilovecake):

plz help

OpenStudy (anonymous):

Find what they both are divisible by. Get to the point were neither number can be divided anymore Both can be divided by 9 18/9=2 27/9=3 2 and 3 have no numbers they can both be simplified to so that is the answer 18/27=2/3

OpenStudy (ilovecake):

thank you!